OpenStudy (anonymous):

Laura must choose a number between 49 and 95 that is a multiple of 3 , 5 , and 9 . Write all the numbers that she could choose. If there is more than one number, separate them with commas.

OpenStudy (rational):

@BSwan

OpenStudy (rational):

sorry, thought this is chinese remainder theorem... but this is very simple lol

OpenStudy (rational):

choose only between 49 and 95

OpenStudy (anonymous):

n=5*9*3 haha

OpenStudy (rational):

9*5 = 45 so they will be of form 45n

OpenStudy (anonymous):

oh lol forgot that 9=3*3